# On the Invariant and Geometric Structure of the Holomorphic Unified Field Theory

### v1

@@ 2 Invariant Geometry of HUFT
Now we take $\kappa$ to denote the Killing form on $\mathfrak{g}$ and $\langle\cdot,\cdot\rangle_{h}$, the fiberwise inner product induced by $h$ on differential forms. We say that gravity and gauge interactions are fully geometrically unified if there exists a single principal HH–bundle $P_{\mathrm{tot}}\to M$, a single connection $\mathcal{A}$ on $P_{\mathrm{tot}}$ with curvature $\mathcal{F}$, and a single $\mathrm{Diff}(M)\times H$–invariant action $S[h,\mathcal{A},\Psi]$ such that the Euler–Lagrange equations are equivalent to the Einstein equations for $h$ coupled to the Yang–Mills equations for $A$ and matter $\Psi$. The unique Bianchi identity $D_{\mathcal{A}}\mathcal{F}=0$ simultaneously yields the Riemann and Yang–Mills Bianchi identities [7, 3]. On $(M,P_{\mathrm{tot}},h,\mathcal{A})$ as above, we define:
@@
where $\langle F,F\rangle_{h}$ is computed using the Killing form $\kappa$ on g and the metric $h$. Then, We wouSS is invariant under $\mathrm{Diff}(M)\times H$. The unique Bianchi identity $D_{\mathcal{A}}\mathcal{F}=0$ splits as $D_{\omega}R=0$ and $D_{A}F=0$. Under an HH–gauge transformation $g:M\to H$:
@@
hence $F\mapsto g^{-1}Fg$ on the internal block. Because the Killing form $\kappa$ on $\mathfrak{g}$ is $\mathrm{Ad}$–invariant, $\mathrm{tr}_{\kappa}(F\wedge*_{h}F)$ is gauge invariant. Each integrand in $S$ is a scalar density built from $h$ (via $_{h}$ and contractions), so the action is also invariant under $\text{Diff}(M)$. The Euler–Lagrange equations of $S$ are:
@@
the Einstein equations for $h$ coupled to the Yang–Mills equations for $\mathcal{A}$ with the same connection $A$ acting on matter. Hence the data $(P_{\mathrm{tot}},\mathcal{A},h)$ satisfy the Definition of full geometric unification.
The Einstein–Hilbert term and $\langle F,F\rangle_{h}$ are $\mathrm{Diff}(M)$–invariant by construction as each integrand is a scalar density obtained from h and F using only natural tensor operations, so the whole action is invariant under pullbacks by diffeomorphisms. Gauge invariance under $H$ holds because $F\mapsto g^{-1}Fg$ and $\kappa$ are $\mathrm{Ad}$–invariant, so $\mathrm{tr}_{\kappa}(F\wedge*_{h}F)$ is gauge invariant. The matter Lagrangian is assumed to be constructed from $h$ and the covariant derivative $D_{\mu}=\partial_{\mu}+\frac{1}{4}\omega^{ab}{}_{\mu}\gamma_{ab}+i\,g\,A_{\mu}^{A}T_{A}$, meaning it is also $\mathrm{Diff}(M)\times H$–invariant this is so that every derivative of a field is replaced by one that transforms covariantly under both spacetime and internal symmetries. Since $\mathfrak{h}=\mathfrak{spin}(1,3)\oplus\mathfrak{g}$, the covariant derivative $D_{\mathcal{A}}$ splits as $D_{\mathcal{A}}=(D_{\omega},D_{A})$ on the two summands. Therefore $(D_{\omega}R,\;D_{A}F)=D_{\mathcal{A}}\mathcal{F}=0$, the Riemann and Yang–Mills Bianchi identities simultaneously. Varying $S$ with respect to $A$ and integrating by parts yields:
@@
the Yang–Mills equations (10). Varying $S$ with respect to $h$ gives:
@@
where $T_{\mu\nu}=T^{\rm YM}_{\mu\nu}+T^{\rm matter}_{\mu\nu}$ and $T^{\rm YM}_{\mu\nu}=\mathrm{tr}_{\kappa}\left(F_{\mu\alpha}F_{\nu}{}^{\alpha}-\frac{1}{4}h_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}\right)$, this implies (9). Since both sectors arise from one variational principle on the same $(h,A)$, the Definition is satisfied.
We let $M_{\mathbb{C}}$ be a complexification of $M$ with complex coordinates $z^{\mu}=x^{\mu}+iy^{\mu}$, and let $g=h+iB$ be a Hermitian tensor field on $M_{\mathbb{C}}$ with $h\in\Gamma(S^{2}T^{*}M)$ and $B\in\Omega^{2}(M,\mathrm{ad}P_{G})$ pulled back from $M$. Now we impose the metric-compatibility condition with the same master connection:
@@
so $h$ is the spacetime metric, and $B$ is an $\mathrm{ad}P_{G}$–valued closed 22–form under $D_{A}$. In the dynamical theory, either by definition or by adding a holomorphic penalty term $\int\sqrt{|h|}\,\langle B-F,B-F\rangle_{h}$ and using the Euler–Lagrange equation, we obtain $B=F$ on the real slice [7, 3], we augment the action by the holomorphic penalty:
@@
where $F=dA+A\wedge A$, $\kappa$ is the Killing form on $\mathfrak{g}$, and $_{h}$ is the Hodge operator determined by $h$. We treat $h,A,B$ as independent fields and we work on an oriented $M$, drop the boundary terms, and use that $_{h}$ is an isomorphism on $2$-forms in four dimensions. Since $S_{\text{pen}}$ is algebraic in $B$:
@@
Because $\delta B$ is arbitrary, the Euler–Lagrange equation from $\delta_{B}S_{\text{pen}}=0$ is:
@@
so the AA–equation of motion contributed by $S_{\text{pen}}$ is:
@@
This means the Euler–Lagrange equation from varying $B$ in (15) enforces the pointwise identification $B=F\quad\text{on }M,$ so the antisymmetric sector coincides with the Yang–Mills curvature on the real slice. The antisymmetric sector of $g$ is the gauge curvature, while all index operations use $h$. Because $S$ is $\mathrm{Diff}(M)\times H$–invariant and built from $(h,A)$, the single Noether identity yields both covariant conservations:
@@
and matter couples by the same covariant derivative $D_{\mu}$ that contains both the spin and internal connections. This shows parallel transport, curvature, symmetries, and dynamics are all governed by the single geometric object $\mathcal{A}$ on $P_{\mathrm{tot}}$, with $h$ supplying measurement such as Hodge duals, index operations. This is precisely full geometric unification in the sense of Definition 1, achieved without inverting the full Hermitian field $g=h+iB$.
In this s is the frame choice on $P_{Spin}$ as $s_{Spin}:M\to P_{Spin}$ picks an orthonormal frame. Pulling back the canonical forms on the bundle gives the coframe and the local Spin connection on M:
@@
The bundle geometry of the bundle $P_{Spin}\times_{M}P_{G}$ as connected by $\mathcal{F,A}$, we have $e^{\alpha}$ as the coframe one form that converts spacetime indices ($\mu,\nu,$...) to local Lorentz indices ($\alpha,\beta,$...). We have a set of one forms on $M$:
@@
The dual to the frame vector fields $e_{\alpha}=e^{\mu}_{\alpha}\partial_{\mu}$ with:
@@
In this $e^{\alpha}$ provides the soldering between the Spin bundle and the the tangent bundle (TM) of the manifold $M$.
@@
with components $X=X^{\mu}(x)\,\partial_{\mu}$. The cotangent bundle $T^{*}M$ has fibers of 1-forms so, the tetrad or coframe $e^{\alpha}=e^{\alpha}{}_{\mu}\,dx^{\mu}$ lives in $\Omega^{1}(M)=\Gamma(T^{*}M)$. The metric $h$ is a section of $S^{2}T^{*}M$ and identifies vectors or covectors through the musical isomorphisms  [^1] Musical isomorphisms have the form: $\sharp=\hat{g}:T^{*}M\to TM,\quad\flat=\hat{g}:TM\to T^{*}M,$ (27) these may as well be referred to as mutually inverse isomorphisms, and the (co)vectors obtained from this way are called metrically equivalent. [42, 43]:
@@
We pick a local orthonormal basis $e_{\alpha}=e_{\alpha}{}^{\mu}\partial_{\mu}$ of $TM$. The coframe $e^{\alpha}=e^{\alpha}{}_{\mu}dx^{\mu}$ is its dual, and:
@@
This is what we mean by the tetrad soldering $P_{\mathrm{Spin}}$ to $TM$: pulling back the solder form on the spin bundle gives $e^{\alpha}$, a $T^{*}M$-valued object.
For $\pi:M_{\mathbb{C}}\to M$ with $z^{\mu}=x^{\mu}+i\,y^{\mu}$, the tangent bundle of the complexification is $TM_{\mathbb{C}}$. An Ehresmann connection on $\pi$ gives, near $y=0$, a splitting:
@@
If $s:M\to P_{Spin}$ is a local selection and $\vartheta^{\alpha}$ is the solder form, then $\vartheta^{\alpha}$ is the $\mathbb{R}^{1,3}$-valued one form on the orthonormal frame or Spin bundle that identifies tangent directions on $M$ with components in a chosen local Lorentz frame. It is the geometric device that connects the principal frame bundle to the base manifold’s tangent bundle.
We let $FM$ be the frame bundle of an $n$-manifold $M$, we have a point $u\in FM$ is a linear isomorphism $u:\mathbb{R}^{n}\to T_{x}M$, a frame at $x=\pi(u)$. The solder form $\vartheta\in\omega^{1}(FM);\mathbb{R}^{n}$ is defined by:
@@
So this means, project a tangent vector $X$ at $u$ down to $T_{x}M$ through $d\pi$, then express it in the frame $u$ by applying $u^{-1}$. It has the property that if $X$ is tangent to the fiber, $d\pi(X)=0$, the $\vartheta(X)=0$. For the right action $R_{g}:FM\to FM,\quad g\in GL(n)$:
@@
On the orthonormal frame bundle $P_{SO(1,3)},$ this holds with $g\in SO(1,3)$.
For frames to the coframe on $e^{\alpha}$ on M, we chose a local section, $s:U\subset M\implies P_{SO(1,3)}$. We pull back the solder form:
@@
These are the tetrad or coframe 1-forms (24).
The relation to the spin bundle is shown when we let $\lambda:P_{Spin}\implies P_{SO(1,3)}$ be the 2-to-1 form covering, where $\lambda$ is the bundle map that realizes the spin structure—the double cover from the Spin principal bundle to the orthonormal frame bundle. We pull $\vartheta$ back along $\lambda$ to get the spin bundle version:
@@
Given the Spin connection $\omega_{\beta}^{\alpha}$ on $P_{Spin}$, the Cartan structure quations are:
@@
for Levi-Civita geometry $T^{\alpha}=0$.
@@
We have the metric identities:
@@
The boundary term is cancelled by the Gibbons–Hawking–York term [30, 31]:
@@
Stationarity $\delta S_{\rm EH}+\delta S_{\rm m}=0$ for arbitrary $\delta h^{\mu\nu}$ gives:
@@
In the Palatini or tetrad formalism [32, 33, 34] one varies $(e^{a},\omega^{ab})$ independently, the torsionless solution reproduces the Levi–Civita case. If $M$ has a boundary, use $S_{\rm GHY}$ and the YM boundary term $\int_{\partial M}\!\langle\delta A\wedge{*}F\rangle$ to enforce Dirichlet data.

### v2

@@ 2 Invariant Geometry of HUFT
We emphasize that adopting $G=SU(5)$ as the minimal-rank unifying group is a statement about geometric compatibility and does not yet by itself resolve known phenomenological constraints of four-dimensional $SU(5)$ GUTs. In particular, integrating out the heavy $X,Y$ gauge bosons generically induces baryon-number violating dimension-six operators of the following schematic form:
$\mathcal{L}_{\Delta B\neq 0}\sim\frac{g_{5}^{2}}{M_{X}^{2}}\,(qqql+\cdots),$ (18)
so experimental limits translate into a lower bound on the effective unification/breaking scale $M_{X}$ and additional selection rules.
Within HUFT, $SU(5)$ should be viewed as the UV internal structure of the holomorphic bundle data, while the real-slice physics is Standard-Model-like after adjoint reduction. A fully realistic model must therefore supplement the present geometric framework with a concrete breaking mechanism and matter-bundle assignment that satisfies proton-decay bounds; we treat these constraints as part of the required phenomenology rather than as automatic consequences of the kinematic packaging.
Now, we take $\kappa$ to denote the Killing form on $\mathfrak{g}$ and $\langle\cdot,\cdot\rangle_{h}$, the fiberwise inner product induced by $h$ on differential forms. {Definition}[Full geometric unification] We say that gravity and gauge interactions are fully geometrically unified if there exists a single principal HH–bundle $P_{\mathrm{tot}}\to M$, a single connection $\mathcal{A}$ on $P_{\mathrm{tot}}$ with curvature $\mathcal{F}$, and a single $\mathrm{Diff}(M)\times H$–invariant action $S[h,\mathcal{A},\Psi]$ such that the Euler–Lagrange equations are equivalent to the Einstein equations for $h$ coupled to the Yang–Mills equations for $A$ and matter $\Psi$. The unique Bianchi identity $D_{\mathcal{A}}\mathcal{F}=0$ simultaneously yields the Riemann and Yang–Mills Bianchi identities Trautman1979 ; WaldGR . On $(M,P_{\mathrm{tot}},h,\mathcal{A})$ as above, we define
@@
where $\langle F,F\rangle_{h}$ is computed using the Killing form $\kappa$ on g and the metric $h$. Then, We will show $S$ is invariant under $\mathrm{Diff}(M)\times H$. The unique Bianchi identity $D_{\mathcal{A}}\mathcal{F}=0$ splits as $D_{\omega}R=0$ and $D_{A}F=0$. Under an HH–gauge transformation $g:M\to H$:
@@
hence $F\mapsto g^{-1}Fg$ on the internal block. Because the Killing form $\kappa$ on $\mathfrak{g}$ is $\mathrm{Ad}$-invariant, $\mathrm{tr}_{\kappa}(F\wedge*_{h}F)$ is gauge-invariant. Each integrand in $S$ is a scalar density built from $h$ (via $_{h}$ and contractions), so the action is also invariant under $\text{Diff}(M)$. The Euler–Lagrange equations of $S$ are
@@
The Einstein equations for $h$ are coupled to the Yang–Mills equations for $\mathcal{A}$ with the same connection $A$ acting on matter. Hence, the data $(P_{\mathrm{tot}},\mathcal{A},h)$ satisfy the Definition of full geometric unification.
The Einstein–Hilbert term and $\langle F,F\rangle_{h}$ are $\mathrm{Diff}(M)$-invariant by construction, as each integrand is a scalar density obtained from h and F using only natural tensor operations, so the whole action is invariant under pullbacks by diffeomorphisms. Gauge invariance under $H$ holds because $F\mapsto g^{-1}Fg$ and $\kappa$ are $\mathrm{Ad}$-invariant, so $\mathrm{tr}_{\kappa}(F\wedge*_{h}F)$ is gauge invariant. The matter Lagrangian is assumed to be constructed from $h$ and the covariant derivative $D_{\mu}=\partial_{\mu}+\frac{1}{4}\omega^{ab}{}_{\mu}\gamma_{ab}+i\,g\,A_{\mu}^{A}T_{A}$, meaning it is also $\mathrm{Diff}(M)\times H$-invariant; this is so that every derivative of a field is replaced by one that transforms covariantly under both spacetime and internal symmetries. Since $\mathfrak{h}=\mathfrak{spin}(1,3)\oplus\mathfrak{g}$, the covariant derivative $D_{\mathcal{A}}$ splits as $D_{\mathcal{A}}=(D_{\omega},D_{A})$ on the two summands, $(D_{\omega}R,\;D_{A}F)=D_{\mathcal{A}}\mathcal{F}=0$, for the Riemann and Yang–Mills Bianchi identities simultaneously. Varying $S$ with respect to $A$ and integrating by parts yields
@@
the Yang–Mills Equation (22). Varying $S$ with respect to $h$ gives
@@
where $T_{\mu\nu}=T^{\rm YM}_{\mu\nu}+T^{\rm matter}_{\mu\nu}$ and $T^{\rm YM}_{\mu\nu}=\mathrm{tr}_{\kappa}\left(F_{\mu\alpha}F_{\nu}{}^{\alpha}-\frac{1}{4}h_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}\right)$; this implies (21). Since both sectors arise from one variational principle on the same $(h,A)$, the Definition is satisfied.
We let $M_{\mathbb{C}}$ be a complexification of $M$ with complex coordinates $z^{\mu}=x^{\mu}+iy^{\mu}$, and let $g=h+iB$ be a Hermitian tensor field on $M_{\mathbb{C}}$ with $h\in\Gamma(S^{2}T^{*}M)$ and $B\in\Omega^{2}(M,\mathrm{ad}P_{G})$ pulled back from $M$.
We use units $c=\hbar=1$ and take local coordinates $x^{\mu}$ to have the length dimension $x^{\mu}]=L$. Then, the Lorentzian metric components are dimensionless, $h_{\mu\nu}]=1$, while a gauge connection one-form $A=A_{\mu}\,dx^{\mu}$ has $A_{\mu}]=L^{-1}$ and its curvature two-form:
$F:=dA+A\wedge A\in\Omega^{2}(M,\mathrm{ad}\,P_{G}),\qquad[F_{\mu\nu}]=L^{-2}.$
Since our antisymmetric sector is identified with curvature on the real slice $B=F$, it is natural to take $B_{\mu\nu}]=L^{-2}$. To form a dimensionless Hermitian packaging field that can be added to $h_{\mu\nu}$, we introduce a fixed length scale $\ell_{\ast}$ equivalently $M_{\ast}:=\ell_{\ast}^{-1}$) and define
$g_{\mu\nu}\;:=\;h_{\mu\nu}\;+\;i\,\ell_{\ast}^{\,2}\,B_{\mu\nu}.$ (25)
On the real slice, where the compatibility mechanism enforces $B=F$, we have $g_{\mu\nu}=h_{\mu\nu}+i\,\ell_{\ast}^{2}F_{\mu\nu}$. But for this paper, we keep it simple and should note that it is a way of packaging the field and should not be read as a literal additive.
Now we impose the metric-compatibility condition with the same master connection:
@@
so $h$ is the spacetime metric, and $B$ is an $\mathrm{ad}P_{G}$-valued closed $2$-form under $D_{A}$. In the dynamical theory, either by definition or by adding a holomorphic penalty term $\int\sqrt{|h|}\,\langle B-F,B-F\rangle_{h}$ and using the Euler–Lagrange equation, we obtain $B=F$ on the real slice Trautman1979 ; WaldGR , we augment the action by the holomorphic penalty:
@@
where $F=dA+A\wedge A$, $\kappa$ is the Killing form on $\mathfrak{g}$, and $_{h}$ is the Hodge operator determined by $h$. A common misunderstanding is that one may simply define the antisymmetric sector of $g$ to be the Yang–Mills curvature. Off shell, however, $F$ is not an independent tensor field; it is constrained to be the curvature of a connection $A$ and therefore satisfies the Bianchi identity $D_{A}F=0$ identically. Introducing an auxiliary adjoint-valued two-form $B$ allows us to keep the Hermitian packaging field $g=h+i\,\ell_{*}^{2}B$ well-defined off the shell. We implement $B=F$ as an equation of motion while preserving both diffeomorphism invariance and gauge invariance, and vary the action cleanly with respect to $(h,A,B)$ without ever needing to invert the full Hermitian tensor $g$. In the large-$\lambda$ limit, $S_{\mathrm{pen}}$ enforces $B\to F$ strongly; for finite $\lambda$, it provides a covariant compatibility mechanism whose Euler–Lagrange equation still sets $B=F$ on the real slice. We let $G$ be compact and reductive with Lie algebra $\mathfrak{g}$ and principal bundle $P_{G}\to M$. Write the adjoint bundle as $\mathrm{ad}\,P_{G}:=P_{G}\times_{\mathrm{Ad}}\mathfrak{g}$. Then,
$B,\;F\in\Omega^{2}(M,\mathrm{ad}\,P_{G}),$
so $B_{\mu\nu}$ and $F_{\mu\nu}$ are $\mathfrak{g}$-valued antisymmetric tensors: in a basis $\{T_{A}\}$ of $\mathfrak{g}$, $B_{\mu\nu}=B_{\mu\nu}^{A}T_{A}$ and $F_{\mu\nu}=F_{\mu\nu}^{A}T_{A}$. Fix an $\mathrm{Ad}$-invariant bilinear form $\kappa:\mathfrak{g}\times\mathfrak{g}\to\mathbb{R}$ such as the Killing form on the semisimple part, or $\kappa(X,Y)=\mathrm{Tr}(\rho(X)\rho(Y))$ in a faithful unitary representation $\rho$. Let $\kappa_{AB}:=\kappa(T_{A},T_{B})$. Now define the $h$-induced inner product on $\mathfrak{g}$-valued two-forms by
$\langle X,Y\rangle_{h}:=\frac{1}{2}\,\kappa_{AB}\,X_{\mu\nu}^{A}\,Y_{\alpha\beta}^{B}\,h^{\mu\alpha}h^{\nu\beta},\qquad X,Y\in\Omega^{2}(M,\mathrm{ad}\,P_{G}),$ (29)
so that $\int_{M}\sqrt{|h|}\,\langle B-F,B-F\rangle_{h}=\int_{M}\kappa\!\left((B-F)\wedge\ast_{h}(B-F)\right)$. Under a gauge transformation $u:M\to G$:
$B\mapsto\mathrm{Ad}_{u^{-1}}B,\qquad F\mapsto\mathrm{Ad}_{u^{-1}}F,$
and $\kappa$ is $\mathrm{Ad}$-invariant, so $\langle B-F,B-F\rangle_{h}$ is gauge invariant. Thus, the identification $B=F$ is an equality in $\Omega^{2}(M,\mathrm{ad}\,P_{G})$; no projection onto a fixed generator is required and such a projection would generally break gauge invariance unless additional adjoint-breaking data are introduced. We treat $h,A,B$ as independent fields and we work on an oriented $M$, drop the boundary terms, and state that $_{h}$ is an isomorphism on $2$-forms in four dimensions. Since $S_{\text{pen}}$ is algebraic in $B$:
@@
Because $\delta B$ is arbitrary, the Euler–Lagrange equation from $\delta_{B}S_{\text{pen}}=0$ is
@@
so the $A$-equation of motion contributed by $S_{\text{pen}}$ is
@@
This means the Euler–Lagrange equation from varying $B$ in (28) enforces the pointwise identification $B=F\quad\text{on }M,$ so the antisymmetric sector coincides with the Yang–Mills curvature on the real slice. The antisymmetric sector of $g$ is the gauge curvature, while all index operations use $h$. Because $S$ is $\mathrm{Diff}(M)\times H$-invariant and built from $(h,A)$, the single Noether identity yields both covariant conservations:
@@
and matter couples by the same covariant derivative $D_{\mu}$ that contains both the spin and internal connections. This shows that parallel transport, curvature, symmetries, and dynamics are all governed by the single geometric object $\mathcal{A}$ on $P_{\mathrm{tot}}$, with $h$ supplying measurements such as Hodge duals and index operations. This is precise full geometric unification in the sense of Definition 2, achieved without inverting the full Hermitian field $g=h+iB$.
In this, s is the frame choice on $P_{Spin}$ as $s_{Spin}:M\to P_{Spin}$ picks an orthonormal frame. Pulling back the canonical forms on the bundle gives the coframe and the local Spin connection on M:
@@
For the bundle geometry of the bundle $P_{Spin}\times_{M}P_{G}$ as connected by $\mathcal{F,A}$, we have $e^{\alpha}$ as the coframe one form that converts spacetime indices ($\mu,\nu,$...) to local Lorentz indices ($\alpha,\beta,$...). We have a set of one forms on $M$ as shown in Figure 1:
@@
The dual to the frame vector fields is $e_{\alpha}=e^{\mu}_{\alpha}\partial_{\mu}$ with
@@
In this, $e^{\alpha}$ provides the soldering between the Spin bundle and the tangent bundle (TM) of the manifold $M$.
@@
with components $X=X^{\mu}(x)\,\partial_{\mu}$. The cotangent bundle $T^{*}M$ has fibers of one-forms, so the tetrad or coframe $e^{\alpha}=e^{\alpha}{}_{\mu}\,dx^{\mu}$ lives in $\Omega^{1}(M)=\Gamma(T^{*}M)$. The metric $h$ is a section of $S^{2}T^{*}M$ and identifies vectors or covectors through the musical isomorphisms. To review, musical isomorphisms have the form:
$\sharp=\hat{g}:T^{*}M\to TM,\quad\flat=\hat{g}:TM\to T^{*}M,$ (41)
these may also be referred to as mutually inverse isomorphisms, and the (co)vectors obtained in this way are called metrically equivalent Music ; Curvature :
@@
We pick a local orthonormal basis $e_{\alpha}=e_{\alpha}{}^{\mu}\partial_{\mu}$ of $TM$. The coframe $e^{\alpha}=e^{\alpha}{}_{\mu}dx^{\mu}$ is its dual, and
@@
This is what we mean by the tetrad soldering $P_{\mathrm{Spin}}$ to $TM$; pulling back the solder form on the spin bundle gives $e^{\alpha}$, a $T^{*}M$-valued object. For $\pi:M_{\mathbb{C}}\to M$ with $z^{\mu}=x^{\mu}+i\,y^{\mu}$, the tangent bundle of the complexification is $TM_{\mathbb{C}}$. An Ehresmann connection on $\pi$ gives, near $y=0$, a splitting:
@@
If $s:M\to P_{Spin}$ is a local selection and $\vartheta^{\alpha}$ is the solder form, then $\vartheta^{\alpha}$ is the $\mathbb{R}^{1,3}$-valued one-form on the orthonormal frame or spin bundle that identifies tangent directions on $M$ with components in a chosen local Lorentz frame. It is the geometric device that connects the principal frame bundle to the base manifold’s tangent bundle.
We let $FM$ be the frame bundle of an $n$-manifold $M$, and we have a point $u\in FM$ which is a linear isomorphism $u:\mathbb{R}^{n}\to T_{x}M$, a frame at $x=\pi(u)$. The solder form $\vartheta\in\omega^{1}(FM);\mathbb{R}^{n}$ is defined by:
@@
So this means that we project a tangent vector $X$ at $u$ down to $T_{x}M$ through $d\pi$, then express it in the frame $u$ by applying $u^{-1}$. It has the property that if $X$ is tangent to the fiber, $d\pi(X)=0$, the $\vartheta(X)=0$. For the right action $R_{g}:FM\to FM,\quad g\in GL(n)$:
@@
On the orthonormal frame bundle $P_{SO(1,3)},$ this holds with $g\in SO(1,3)$. For frames to the coframe on $e^{\alpha}$ on M, we chose a local section, $s:U\subset M\implies P_{SO(1,3)}$. We pull back the solder form:
@@
These are the tetrad or coframe one-forms (38). The relation to the spin bundle is shown when we let $\lambda:P_{Spin}\implies P_{SO(1,3)}$ be the 2-to-1 covering, where $\lambda$ is the bundle map that realizes the spin structure—the double cover from the Spin principal bundle to the orthonormal frame bundle. We pull $\vartheta$ back along $\lambda$ to get the spin bundle version:
@@
Given the Spin connection $\omega_{\beta}^{\alpha}$ on $P_{Spin}$, the Cartan structure equations are:
@@
for Levi–Civita geometry $T^{\alpha}=0$. For completeness, we recall why the metric, Levi–Civita derivation used below is equivalent to the first-order tetrad or Palatini derivation when torsion vanishes. Given a coframe $e^{a}=e^{a}{}_{\mu}\,dx^{\mu}$ with inverse $e_{a}{}^{\mu}$, the Lorentzian metric is:
@@
Variations are related by:
[…]
In the first-order Hilbert–Palatini formulation, the gravitational action can be written as follows:
$S_{\rm HP}[e,\omega]=\frac{1}{4\kappa}\int_{M}\epsilon_{abcd}\,e^{a}\wedge e^{b}\wedge R^{cd}(\omega),$ (53)
where $R^{ab}(\omega)=d\omega^{ab}+\omega^{a}{}_{c}\wedge\omega^{cb}$ and $\epsilon_{abcd}$ is the Levi–Civita symbol in the local Lorentz frame. Varying independently in $(e^{a},\omega^{ab})$ yields $\delta_{\omega}S_{\rm HP}=0\ \Rightarrow\ T^{a}:=de^{a}+\omega^{a}{}_{b}\wedge e^{b}=0$, so the unique solution is the torsionless spin connection $\omega=\omega(e)$, inserting $\omega(e)$ into $\delta_{e}S_{\rm HP}=0$ gives the Einstein equation, which is equivalent to the metric variation of the Einstein–Hilbert action $S_{\rm EH}[h]=\frac{1}{2\kappa}\int_{M}\sqrt{|h|}\,R(h)$.
In the torsionless sector that is relevant here, the tetrad or Palatini and metric or Levi–Civita formulations produce the same equations of motion and the same solution space modulo local Lorentz gauge transformations of the tetrad DadhichPons2012 .
[…]
@@
The boundary term is cancelled by the Gibbons–Hawking–York term York1972 ; GibbonsHawking1977 :
@@
Stationarity $\delta S_{\rm EH}+\delta S_{\rm m}=0$ for arbitrary $\delta h^{\mu\nu}$ gives
@@
In the Palatini or tetrad formalism Palatini1919 ; HehlEtAl1976 ; Holst1996 one varies $(e^{a},\omega^{ab})$ independently, and the torsionless solution reproduces the Levi–Civita case. If $M$ has a boundary, use $S_{\rm GHY}$ and the YM boundary term $\int_{\partial M}\!\langle\delta A\wedge{*}F\rangle$ to enforce Dirichlet data.